A box contains 2 one-dollar bills, 1 five-dollar bill, and 1 ten-dollar bill. A player blindly and randomly draws bills one at a time without replacement from the box until a ten-dollar bill is drawn. Then the game stops. All bills drawn are kept by the player. What is the probability of winning exactly $16?
Answer by edjones(8007)

A $1 bill must remain in the box to have exactly $16 and the last bill taken must be the $10.
Probability that the last bill is a $1 on a random draw is 2/4=1/2
Of the 1/2 when this occurs the probability that the third bill taken is the $10 is 1/3.
1/2 * 1/3 = 1/6 the probability of winning exactly $16.
.
Ed